# Extreme mass-ratio inspiral within an ultralight scalar cloud I. Scalar radiation

### v1

@@ I Introduction
Figure 2: The infinity and horizon energy fluxes $\dot{E}^{s,\infty/H}$ associated with the scalar radiation $\Phi^{(1,1)}$ of the $(\ell_{c}=m_{c}=1,n_{c}=0)$ complex scalar cloud around a Kerr BH of spin $a=0.88M$, where the complex scalar field has mass $\mu M=0.3$. The blue lines represent the infinity flux $\dot{E}^{s,\infty}$, while the orange lines show the horizon flux $\dot{E}^{s,H}$ but with an overall minus sign, given that $\dot{E}^{s,H}$ is always negative here. The solid and dashed lines are the fluxes obtained via the semi-analytical approach in this work and the numerical approach in [105], respectively. The green solid and dashed lines show the relative fractional differences in percentages between these two approaches in the horizon and infinity fluxes, respectively. The red vertical line marks the threshold radius $r_{0}$ where the infinity flux has a sharp decrease.
@@
As a first step, we extend the MTF to EMRIs around “dirty” BHs in astrophysical environments. Using this extension, we compute the scalar radiation emitted by a complex scalar cloud perturbed by a stellar-mass object in a circular and equatorial orbit around a rotating supermassive BH, evaluate the associated energy and angular momentum fluxes, and compare our results with those of [105]. Our semi-analytical approach, though independent from the more numerical methods in [105], yields consistent results for the scalar radiation profile and fluxes at the horizon and infinity (i.e., see Fig. 2 for a comparison of the fluxes) and is more efficient given the analytical simplifications made [^1] We have also corrected several minor errors in [105], including a missing normalization factor in their horizon flux.. In addition to the dipolar clouds studied in [105], we also extend this analysis to quadrupolar clouds and show that different cloud configurations can affect the secondary’s inspiral differently. For example, the quadrupolar cloud’s scalar radiation accelerates the secondary’s inspiral at small orbital radii, whereas the dipolar cloud slows it down. This work establishes key techniques for our forthcoming study on gravitational radiation and provides essential scalar data for constructing the source term in the modified Teukolsky equation.
@@ II EMRI within an ultralight scalar cloud
A complex scalar field with mass $\mu$ and minimally coupled to gravity in GR is described by the action [46]
$S=\int d^{4}x\sqrt{-g}\left(\frac{R}{16\pi}-\partial_{\mu}\Phi^{*}\partial^{\mu}\Phi-\mu^{2}\Phi^{*}\Phi+\mathcal{L}_{\mathrm{m}}\right)\,,$ (1)
where $\mathcal{L}_{\mathrm{m}}$ is the Lagrangian density of additional, non-minimally coupled matter fields. The corresponding equations of motion are given by
@@
$T_{\mu\nu}^{\Phi}=2\partial_{(\mu}\Phi\partial_{\nu)}\Phi^{*}-g_{\mu\nu}\left(\partial_{\alpha}\Phi\partial^{\alpha}\Phi^{*}+\mu^{2}\Phi^{*}\Phi\right)\,,$ (4)
@@ III Evaluation of the scalar equation
Figure 4: The radial source $\mathscr{S}_{\ell m}(r)$ at $\ell=2,3$ and $r_{0}=20M$ for the $(\ell_{c}=m_{c}=1,n_{c}=0)$ scalar cloud around a Kerr BH of spin $a=0.88M$ and with scalar mass $\mu M=0.3$. As discussed in Sec. III, the sources with $\ell$ and $m$ of opposite parity are zero, so they are not displayed here. We also only include positive $m$ for simplicity. The solid and dashed lines are the sources obtained via the semi-analytical approach in this work and the numerical approach in [105], respectively. The dotted lines show the relative fractional differences in percentages between these two approaches when computing $\left|\mathscr{S}_{\ell m}(r)\right|$.
@@
where $\Delta(r)=r^{2}-2Mr+a^{2}$, $\Gamma(r,\theta)=r+ia\cos{\theta}$, and an overhead bar stands for complex conjugation. After projecting $\mathcal{S}_{\Phi}^{(1,1)}$ onto $e_{a}^{\mu}$, one finds
@@
Figure 5: The radial source $\mathscr{S}_{\ell m}(r)$ for the same cloud configuration in Fig. 4 at $\ell=6,7$ and $r_{0}=20M$ for positive $m$. The solid and dashed lines are the sources obtained via the semi-analytical approach in this work and the numerical approach in [105], respectively. The dotted lines show the relative fractional differences in percentages between these two approaches in the values of $\left|\mathscr{S}_{\ell m}(r)\right|$.
@@ V Energy and angular momentum fluxes
$j_{\mu}^{\Phi}=-i\left(\Phi^{*}\partial_{\mu}\Phi-\Phi\partial_{\mu}\Phi^{*}\right)\,,$ (56)
@@
Since the angular momentum flux is completely determined by the energy flux in Eq. (61c) for the circular orbits considered in this work, we will focus on the energy flux for the rest of this work. Inserting the ${}_{0}\mathscr{R}_{\ell m\omega}^{\Phi}(r)$ found in Sec. IV into Eq. (61), we obtain the horizon and infinity fluxes for a range of $r_{0}\in[4.1,50.1]M$ for the $\ell_{c}=m_{c}=1$ cloud. For a detailed comparison with the results in [105], the infinity flux of each $(\ell,m)$ mode up to $\ell=10$ is shown in Fig. 9 for a secondary on a prograde orbit at $r_{0}=20M$ within the $\ell_{c}=m_{c}=1$ cloud. Note that Fig. 6 of [105] only pulls out a factor of $\epsilon^{2}M_{c}/M$, while we have pulled out a factor of $\epsilon^{2}\zeta^{2}$ in Fig. 9, so a factor of $(\mu M)^{6}$ needs to be removed from the former for comparison purposes. The results found via our semi-analytical approach agree with those obtained from the more numerical approach of [105], with a relative fractional difference of $<5\%$ for all the modes present in Fig. 9, except at $\ell=8,m=2$, where the relative fractional difference is $\sim 11.4\%$. This small discrepancy may arise from differences in constructing the radial source $\mathscr{S}_{\ell m}(r)$ in Eq. (50) (our semi-analytical approach versus the numerical approach in [105]), the precise boundary conditions and cutoff locations near the horizon and infinity used when calculating the homogeneous solutions ${}_{0}\mathscr{R}_{\ell m\omega}^{\textrm{up},\textrm{in}}(r)$, and the slightly different spin value used in [105] ($a=0.877M$ versus $a=0.88M$ here. When the flux becomes very small at $\ell=8,m=2$, higher precision is required for the source term and the homogeneous solutions when constructing Green’s function to calculate the flux accurately, so the discrepancy between these two approaches becomes larger. However, since when $\ell=8,m=2$, the infinity flux is already about $10^{12}$ times smaller than the dominant contribution at $\ell=m=3$, this discrepancy is negligible. The mode-by-mode contribution to the infinity flux peaks at $\ell=m=3$ because the quadrupolar mode ($\ell_{g}=m_{g}=2$) of GWs driven by the secondary has the largest amplitude near spatial infinity. The same feature also appears for the $\ell_{c}=m_{c}=2$ cloud, where the infinity flux peaks at $\ell=m=4$, as shown in Fig. 10.
@@ VI Discussion
The source term in the scalar sector consists of terms from the background scalar cloud profile $\Phi^{(1,0)}$ and the perturbed metric $h_{\mu\nu}^{(0,1)}$, which is driven by the perturbations of the secondary body on the central BH. For the calculation of $\Phi^{(1,0)}$, we employ Leaver’s method [130, 131]. For the comptuation of $h_{\mu\nu}^{(0,1)}$, we directly use the Lorenz-gauge reconstruction method and the associated Mathematica notebooks developed by [96, 98, 99]. To simplify the source terms analytically, we project the source term to the NP basis and replace all the NP directional derivatives and spin coefficients with Chandrasekhar operators, which satisfy certain commutation relations. To convert the source term into an explicitly separable form, we further employ the method of [141] to decompose certain radial-angular factors into Fourier series of the polar angle. We then project this explicitly-separable source onto the spin-00 spheroidal harmonics to extract its radial part. Our semi-analytical treatment of the source term agrees very well with the results obtained through a much more numerical approach [105]. Our procedure to compute the source term in the scalar sector can be directly applied to the gravitational sector. Due to the more complicated structure of the source term within the modified Teukolsky equation, this semi-analytical approach might prove more efficient and accurate when calculating the gravitational flux.
@@ VII Acknowledgements
We are grateful to Conor Dyson and Thomas F.M. Spieksma for sharing the data and assisting with the verification of calculations in [105]. We thank Richard Brito and Yanbei Chen for helpful discussions related to this project. We also thank Sam Dolan for providing the Kerr-Lorenz-Circ Mathematica code developed in [96, 98, 99] and clarifying its usage. D. L. and N. Y. acknowledge support from the Simons Foundation through Award No. 896696, NSF Grant No. PHY-2207650, and the National Aeronautics and Space Administration through award 80NSSC22K0806. P. B. acknowledges support from the Dutch Research Council (NWO) with file number OCENW.M.21.119. C. W.’s research is supported by the Simons Foundation (Award No. 568762), the Brinson Foundation, and the National Science Foundation (via Grants No. PHY-2011961 and No. PHY-2011968). This work makes use of the Kerr-Lorenz-Circ Mathematica code developed by Sam Dolan [96, 98, 99] and the Black Hole Perturbation Toolkit [134]. Our calculations use the Illinois Campus Cluster, a computing resource that is operated by the Illinois Campus Cluster Program (ICCP) in conjunction with the National Center for Supercomputing Applications (NCSA), and is supported by funds from the University of Illinois Urbana-Champaign (UIUC). Some calculations were also conducted in the Resnick High Performance Computing Center, a facility supported by the Resnick Sustainability Institute at the California Institute of Technology.
@@ Appendix B Complete form of the source term $\mathcal{S}_{\Phi}^{(1,1)}$
where we have dropped the arguments and the subscripts labeling the modes of all the functions for simplicity. Recall that ${}_{s}R^{\tilde{a}\tilde{b}}\equiv{}_{s}R^{\tilde{a}\tilde{b}}_{\ell_{g}m_{g}\omega_{g}}(r)$ and ${}_{s}S\equiv{}_{s}Y_{\ell_{g}m_{g}}(\theta)$ are the radial and angular parts of the reconstructed metric $h_{\tilde{a}\tilde{b}}^{(0,1)}$ defined in Eq. (28), respectively, where the indices $\tilde{a},\tilde{b}$ label the components in the unnormalized Kinnersley tetrad $\tilde{e}_{a}^{\mu}=\{l_{+}^{\mu},l_{-}^{\mu},m_{+}^{\mu},m_{-}^{\mu}\}$ defined in Eq. (III). The function ${}_{0}R^{h}\equiv{}_{0}R^{h}_{\ell_{g}m_{g}\omega_{g}}(r)$ is the radial part of the trace $h$ of the reconstructed metric $h_{\tilde{a}\tilde{b}}^{(0,1)}$.

### v2

@@ I Introduction
Figure 2: The infinity and horizon energy fluxes $\dot{E}^{s,\infty/H}$ (left axis) associated with the scalar radiation $\Phi^{(1,1)}$ of the $(\ell_{c}=m_{c}=1,n_{c}=0)$ complex scalar cloud around a Kerr BH of spin $a=0.88M$, where the complex scalar field has mass $\mu M=0.3$. The blue lines represent the infinity flux $\dot{E}^{s,\infty}$, while the orange lines show the horizon flux $\dot{E}^{s,H}$ but with an overall minus sign, given that $\dot{E}^{s,H}$ is always negative here. The solid and dashed lines are the fluxes obtained via the approach in this work and the one in [105], respectively. The green solid and dashed lines show the relative fractional differences in percentages (right axis) between these two approaches in the horizon and infinity fluxes, respectively. The red vertical line marks the threshold radius $r_{0}$ where the infinity flux has a sharp decrease.
@@
As a first step, we extend the MTF to EMRIs around “dirty” BHs in astrophysical environments. Using this extension, we compute the scalar radiation emitted by a complex scalar cloud perturbed by a stellar-mass object in a circular and equatorial orbit around a rotating supermassive BH, evaluate the associated energy and angular momentum fluxes, and compare our results with those of [105]. Our semi-analytical approach, though independent from the more numerical methods in [105], yields consistent results for the scalar radiation profile and fluxes at the horizon and infinity (i.e., see Fig. 2 for a comparison of the fluxes) and is more efficient given the analytical simplifications made. [^1] We have also corrected two minor errors in [105]: the spheroidal separation constant of the scalar field equation and the normalization factor for the horizon flux. In addition to the dipolar clouds studied in [105], we also extend this analysis to quadrupolar clouds and show that different cloud configurations can affect the secondary’s inspiral differently. For example, the quadrupolar cloud’s scalar radiation accelerates the secondary’s inspiral at small orbital radii, whereas the dipolar cloud slows it down. This work establishes key techniques for our forthcoming study on gravitational radiation and provides essential scalar data for constructing the source term in the modified Teukolsky equation.
@@ II EMRI within an ultralight scalar cloud
A complex scalar field $\Phi$ with mass $\mu$ and minimally coupled to gravity in GR is described by the action [46]
$S=\int d^{4}x\sqrt{-g}\left(\frac{R}{16\pi}-\partial^{\mu}\Phi\partial_{\mu}\bar{\Phi}-\mu^{2}\Phi\bar{\Phi}+\mathcal{L}_{\mathrm{m}}\right)\,,$ (1)
where $\mathcal{L}_{\mathrm{m}}$ is the Lagrangian density of additional, non-minimally coupled matter fields, and an overhead bar stands for complex conjugation. The corresponding equations of motion are given by
@@
$T_{\mu\nu}^{\Phi}=2\partial_{(\mu}\Phi\partial_{\nu)}\bar{\Phi}-g_{\mu\nu}\left(\partial_{\alpha}\Phi\partial^{\alpha}\bar{\Phi}+\mu^{2}\Phi\bar{\Phi}\right)\,,$ (4)
@@ III Evaluation of the scalar equation
Figure 4: The radial source $\mathscr{S}_{\ell m}(r)$ (left axis) at $\ell=2,3$ and $r_{0}=20M$ for the $(\ell_{c}=m_{c}=1,n_{c}=0)$ scalar cloud around a Kerr BH of spin $a=0.88M$ and with scalar mass $\mu M=0.3$. As discussed in Sec. III, the sources with $\ell$ and $m$ of opposite parity are zero, so they are not displayed here. We also only include positive $m$ for simplicity. The solid and dashed lines are the sources obtained via the approach in this work and the one in [105], respectively. The dotted lines show the relative fractional differences in percentages (right axis) between these two approaches when computing $\left|\mathscr{S}_{\ell m}(r)\right|$.
@@
where $\Delta(r)=r^{2}-2Mr+a^{2}$ and $\Gamma(r,\theta)=r+ia\cos{\theta}$. After projecting $\mathcal{S}_{\Phi}^{(1,1)}$ onto $e_{a}^{\mu}$, one finds
@@
Figure 5: The radial source $\mathscr{S}_{\ell m}(r)$ (left axis) for the same cloud configuration in Fig. 4 at $\ell=6,7$ and $r_{0}=20M$ for positive $m$. The solid and dashed lines are the sources obtained via the approach in this work and the one in [105], respectively. The dotted lines show the relative fractional differences in percentages (right axis) between these two approaches in the values of $\left|\mathscr{S}_{\ell m}(r)\right|$.
@@ V Energy and angular momentum fluxes
$j_{\mu}^{\Phi}=-i\left(\bar{\Phi}\partial_{\mu}\Phi-\Phi\partial_{\mu}\bar{\Phi}\right)\,,$ (56)
@@
Since the angular momentum flux is completely determined by the energy flux in Eq. (61c) for the circular orbits considered in this work, we will focus on the energy flux for the rest of this work. Inserting the ${}_{0}\mathscr{R}_{\ell m\omega}^{\Phi}(r)$ found in Sec. IV into Eq. (61), we obtain the horizon and infinity fluxes for a range of $r_{0}\in[4.1,50.1]M$ for the $\ell_{c}=m_{c}=1$ cloud. For a detailed comparison with the results in [105], the infinity flux of each $(\ell,m)$ mode up to $\ell=10$ is shown in Fig. 9 for a secondary on a prograde orbit at $r_{0}=20M$ within the $\ell_{c}=m_{c}=1$ cloud. Note that Fig. 6 of [105] only pulls out a factor of $\epsilon^{2}M_{c}/M$, while we have pulled out a factor of $\epsilon^{2}\zeta^{2}$ in Fig. 9, so a factor of $(\mu M)^{6}$ needs to be removed from the former for comparison purposes. The results found via our approach agree with those obtained from the approach of [105], with a relative fractional difference of $<5\%$ for all the modes present in Fig. 9, except at $\ell=8,m=2$, where the relative fractional difference is $\sim 11.4\%$. This small discrepancy may arise from differences in constructing the radial source $\mathscr{S}_{\ell m}(r)$ in Eq. (50) (our approach versus the approach in [105]), the precise boundary conditions and cutoff locations near the horizon and infinity used when calculating the homogeneous solutions ${}_{0}\mathscr{R}_{\ell m\omega}^{\textrm{up},\textrm{in}}(r)$, and the slightly different spin value used in [105] ($a=0.877M$ versus $a=0.88M$ here. When the flux becomes very small at $\ell=8,m=2$, higher precision is required for the source term and the homogeneous solutions when constructing Green’s function to calculate the flux accurately, so the discrepancy between these two approaches becomes larger. However, since when $\ell=8,m=2$, the infinity flux is already about $10^{12}$ times smaller than the dominant contribution at $\ell=m=3$, this discrepancy is negligible. The mode-by-mode contribution to the infinity flux peaks at $\ell=m=3$ because the quadrupolar mode ($\ell_{g}=m_{g}=2$) of GWs driven by the secondary has the largest amplitude near spatial infinity. The same feature also appears for the $\ell_{c}=m_{c}=2$ cloud, where the infinity flux peaks at $\ell=m=4$, as shown in Fig. 10.
@@ VI Discussion
The source term in the scalar sector consists of terms from the background scalar cloud profile $\Phi^{(1,0)}$ and the perturbed metric $h_{\mu\nu}^{(0,1)}$, which is driven by the perturbations of the secondary body on the central BH. For the calculation of $\Phi^{(1,0)}$, we employ Leaver’s method [130, 131]. For the comptuation of $h_{\mu\nu}^{(0,1)}$, we directly use the Lorenz-gauge reconstruction method and the associated Mathematica notebooks developed by [96, 98, 99]. To simplify the source terms analytically, we project the source term to the NP basis and replace all the NP directional derivatives and spin coefficients with Chandrasekhar operators, which satisfy certain commutation relations. To convert the source term into an explicitly separable form, we further employ the method of [141] to decompose certain radial-angular factors into Fourier series of the polar angle. We then project this explicitly-separable source onto the spin-00 spheroidal harmonics to extract its radial part. Our semi-analytical treatment of the source term agrees very well with the results obtained through a more numerical approach [105]. Our procedure to compute the source term in the scalar sector can be directly applied to the gravitational sector. Due to the more complicated structure of the source term within the modified Teukolsky equation, this semi-analytical approach might prove more efficient and accurate when calculating the gravitational flux.
@@ VII Acknowledgements
We are grateful to Conor Dyson and Thomas F.M. Spieksma for sharing the data and assisting with the verification of calculations in [105]. We thank Richard Brito and Yanbei Chen for helpful discussions related to this project. We also thank Sam Dolan for providing the Kerr-Lorenz-Circ Mathematica code developed in [96, 98, 99] and clarifying its usage. We appreciate the insightful comments on the final draft of this manuscript from Conor Dyson, Thomas F.M. Spieksma, and Richard Brito. D. L. and N. Y. acknowledge support from the Simons Foundation through Award No. 896696, NSF Grant No. PHY-2207650, and the National Aeronautics and Space Administration through award 80NSSC22K0806. P. B. acknowledges support from the Dutch Research Council (NWO) with file number OCENW.M.21.119. C. W.’s research is supported by the Simons Foundation (Award No. 568762), the Brinson Foundation, and the National Science Foundation (via Grants No. PHY-2011961 and No. PHY-2011968). This work makes use of the Kerr-Lorenz-Circ Mathematica code developed by Sam Dolan [96, 98, 99] and the Black Hole Perturbation Toolkit [134]. Our calculations use the Illinois Campus Cluster, a computing resource that is operated by the Illinois Campus Cluster Program (ICCP) in conjunction with the National Center for Supercomputing Applications (NCSA), and is supported by funds from the University of Illinois Urbana-Champaign (UIUC). Some calculations were also conducted in the Resnick High Performance Computing Center, a facility supported by the Resnick Sustainability Institute at the California Institute of Technology.
@@ Appendix B Complete form of the source term $\mathcal{S}_{\Phi}^{(1,1)}$
where we have dropped the arguments and the subscripts labeling the modes of all the functions for simplicity. Recall that ${}_{s}R^{\tilde{a}\tilde{b}}\equiv{}_{s}R^{\tilde{a}\tilde{b}}_{\ell_{g}m_{g}\omega_{g}}(r)$ and ${}_{s}Y\equiv{}_{s}Y_{\ell_{g}m_{g}}(\theta)$ are the radial and angular parts of the reconstructed metric $h_{\tilde{a}\tilde{b}}^{(0,1)}$ defined in Eq. (28), respectively, where the indices $\tilde{a},\tilde{b}$ label the components in the unnormalized Kinnersley tetrad $\tilde{e}_{a}^{\mu}=\{l_{+}^{\mu},l_{-}^{\mu},m_{+}^{\mu},m_{-}^{\mu}\}$ defined in Eq. (III). The function ${}_{0}R^{h}\equiv{}_{0}R^{h}_{\ell_{g}m_{g}\omega_{g}}(r)$ is the radial part of the trace $h$ of the reconstructed metric $h_{\tilde{a}\tilde{b}}^{(0,1)}$.